# On a mean-field Pontryagin minimum principle for stochastic optimal control

### v1

@@ 7. Conclusions
A deterministic mean-field formulation of the PMP has been presented. Similar to the well-known forward-backward SDE formulation of the stochastic PMP [2] and its numerical implementation; see e.g. [5], a regression problem (27) arises when discretized as an interacting particle system. Several numerical implementations have been discussed. Contrary to its stochastic counterpart, the mean-field approach leads to smooth particle trajectories which should help with reducing the variance in the involved approximations. Furthermore, the freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively. Furthermore, a forward simulation of the coupled Hamiltonian ODEs is possible in case of infinite horizon optimal control problems. Numerical results have been presented for a controlled inverted pendulum as well as a controlled Lorenz-63 system. While the numerical experiments focus on infinite horizon control problems and the Schrödinger bridge approximation, application of the proposed methodology to NMPC [23, 7] and variational approximations will be investigated in future work. Another area of interest is provided by extensions to partially observed optimal control problems [22, 1, 17, 29, 24].

### v3

@@ 8. Conclusions
A deterministic mean-field formulation of the PMP has been presented. Similar to the well-known forward-backward SDE formulation of the stochastic PMP [5] and its numerical implementation, a regression problem (27) arises when discretized as an interacting particle system. Several numerical implementations have been discussed. Contrary to its stochastic counterpart, the McKean–Pontryagin minimum principle leads to smooth particle trajectories which should help with reducing the variance in the involved approximations. The freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively. Furthermore, a forward simulation of the coupled Hamiltonian ODEs is possible in case of infinite horizon optimal control problems. Numerical results have been presented for a controlled inverted pendulum as well as a controlled Lorenz-63 system. An application of the proposed methodology to NMPC [30, 11] and a variational approximation of the McKean–Pontryagin formulation has also been investigated numerically. An important area of future work is provided by extensions to partially observed optimal control problems [29, 1, 23, 36, 31]. Here a combination of the control formulation (77) with the ensemble Kalman filter [4] appears as a particularly promising direction in the context of NMPC.